OpenStudy (anonymous):

My final is worth 25% of my grade. I have a 79 now. What is the minimum score I can make to get a B in the class? And an A?

OpenStudy (anonymous):

is this a real question or is this your grade?? Either way: if 75% of you grade is a 79 and you still have 25% to complete 75% of 79 + 25% of Final = Grade so if you want a B, which I believe is a 83: 0.75*79+0.25*F = 83; 0.25*F = 23.75 F = 95; if you want an A, which is a 93 (I think) 0.75*79+0.25*F = 93 F = 135; so to get a B you need to get a 95 on your final, and to get an A in the class you need a 135.